# Zeyu Shen1    Basileal Imana1    Tong Wu1    Chong Xiang2    Prateek Mittal1    Aleksandra Korolova1

### v1

@@ 3.2 Robustness Analysis
Performance of Algorithm 1 depends on both NLI and MIS, and thus, of course, the robustness of our proposed system will depend on how NLI treats malicious documents. We assume that the NLI model has an error probability of at most $\epsilon_{1}$ when comparing two benign answers (i.e., it incorrectly labels them as contradictory), and an error probability of at most $\epsilon_{2}$ when comparing a benign answer and a malicious answer (i.e., it incorrectly fails to detect contradiction). Formally, for each pair of answers $(y_{i},y_{j})$ produced from documents $(x_{i},x_{j})$: if both $x_{i}$ and $x_{j}$ are benign, then $\text{NLI}(y_{i},y_{j})$ outputs “non-contradictory” with probability at least $1-\epsilon_{1}$; if exactly one of $x_{i},x_{j}$ is malicious, then $\text{NLI}(y_{i},y_{j})$ outputs “contradictory” with probability at least $1-\epsilon_{2}$. We place larger tolerance on $\epsilon_{2}$ because adversaries may craft malicious documents in such a way that induces larger NLI error rates.
We make no assumption on NLI output when both are malicious.
Theorem 3.1.
@@
The assumption that NLI is able to find contradictions over LLM’s isolated answers is justified by the targeted attack we consider in our threat model, where an adversary aims to manipulate the model’s output to induce a specific outcome, such as inclusion of their own product in Google AI Overview. Thus, for the attack to be meaningful in this setting, the malicious document should diverge from the information in benign documents, promoting an alternative product that would not otherwise appear in the output [^4] We acknowledge the possibility that this assumption may not hold in reality, and that a sophisticated adversary may design adaptive attacks based off from this. We discuss this possibility in more detail in Appendix D.2.. The assumption of bounded corruption stems from the practical difficulty of manipulating many top-ranked documents in RAG-based search systems that rely on the strength of modern information retrieval systems against SEO [26, 1, 56].
For the NLI model, we use DeBERTa-v3-large-mnli-fever-anli-ling-wanli [30] checkpoint, which achieves state-of-the-art 91.2% / 90.8% accuracy on the MNLI-m/mm test splits (benign) and 70.2% on the ANLI test set (adversarial). These numbers indicate both high everyday reliability and strong robustness to deliberately hard contradictions.
Theorem 3.1 (proved in Appendix B.1) provides theoretical validation that ReliabilityRAG via MIS is provably robust since the probability of selecting a malicious document vanishes as the number of retrieved documents $k$ grows large, provided the NLI error rates $\epsilon_{1},\epsilon_{2}$ and the number of malicious documents $k^{\prime}$ satisfy the specified conditions. To better understand the practical robustness implications in small-$k$ regimes, we present empirical results demonstrating practical robustness in such regimes in Appendix B.1.1. In addition, in Appendix B.2, we show that when $\epsilon_{1}=\epsilon_{2}=0$ (perfect NLI), the MIS is exactly the set of benign documents when $k^{\prime}<\frac{k}{2}$, i.e., we guarantee perfect robustness and utility.
@@ 4 Cardinal-Reliability Setting: Weighted Sample and Aggregate Framework
In addition, thus far, we have primarily focused on the ordinal-reliability setting. However, many retrieval systems provide, in addition to a rank, a non-negative reliability score $r(x_{i})$ for each retrieved document $x_{i}$ [13]. We can naturally convert these scores into normalized weights by setting $w(x_{i})=\frac{r(x_{i})}{\sum_{j=1}^{k}r(x_{j})}$, so a larger weight signifies greater trustworthiness. Cardinal reliability therefore makes the distribution of trust explicit: In the search scenario, depending on the query, one could encounter various weight distributions. For example, for queries with many reliable sources, such as encyclopedia-type queries, the weights among the top $k$ retrieved documents may be rather uniform; on the other hand, for niche queries, the highest ranked documents may carry high weights with a sharp drop-off for the lower ranked ones.
@@ 4.1 Weighted Sample and Aggregate Framework
Instantiation of $\mathcal{A}$ with MIS-based Document Selection.
We can instantiate the aggregator $\mathcal{A}$ with MIS-based document selection (Stage 2 and 3 of Algorithm 1) by sending $a_{t}$’s as the isolated answers and contexts $S_{t}$’s as the documents. The ranking over the contexts in the instantiation is defined based on the rankings of the documents inside them. Since each context $\mathcal{S}_{t}$ is a tuple of documents, we can rank them lexicographically based on the ranks of the documents they contain. [^6] Again, this is one of many possible design choices. With this instantiation, we are able to control the running time of MIS via the choice of the number of sampling rounds $T$, as the MIS is now computed on a graph with $T$ vertices instead of $k$.
@@ 5.1 Experimental Setup
Datasets.
We evaluate on three open-domain QA datasets: RealtimeQA (RQA) [28], Natural-Questions (NQ) [32], TriviaQA (TQA) [27], and a long-form Biography generation dataset (Bio) [31]. These datasets have been widely used to study the accuracy and robustness of RAG systems [61, 60, 58], making them well-suited benchmarks for evaluating our method. The detailed setup for datasets is presented in Appendix C.1. Due to limited space, we present results for TQA in Appendix C.2, which largely mirror the results for RQA and NQ.
LLMs and RAG Settings.
We run experiments using three LLMs as the generators in our RAG pipelines: Mistral-7B-Instruct-v0.2 [24], Llama3.2-3B-Instruct [40], and GPT-4o-mini [50]. We set temperature to 0 for all experiments. When testing MIS, we use the top $k=10$ passages. For Sampling + MIS, we use the top $k=50$ documents, since one major motivation for the weighted sample and aggregate framework is scalability. We set context size $m=2$ and number of sampling rounds $T=20$. For the weights, we use the exponentially decaying weights and set $w(x_{i})\propto\gamma^{i-1}$, where $\gamma=0.9$. We present detailed ablation studies and discussions on the choice of parameters and their impact in Appendix C.4.
Evaluation Scenarios.
We assess our algorithm’s performance and robustness under two distinct scenarios: a benign setting where no adversarial attack is performed and an adversarial attack setting. We consider the benign setting as defenses risk filtering out helpful content along with harmful attacks, reducing accuracy even when there is no attack [61]. For the adversarial setting, we simulate targeted corruption of one document at specific ranks: positions 1 (highest) and 10 (lowest) for $k=10$, and positions 1 (highest), 25 (middle), and 50 (lowest) for $k=50$. We also evaluate on multi-position attacks.
@@
Evaluation Metrics.
We evaluate all methods based on their ability to generate accurate responses, both in benign conditions and under attack. We report GPT-4o-judged answer-accuracy for QA datasets and a GPT-4o judge score (0–100) for Bio. Due to limited space, we evaluate Attack Success Rate (ASR) in Appendix C.3.1.
Baselines.
We compare our proposed reliability-aware methods against several baselines: Vanilla RAG, which concatenates all retrieved passages without any defense mechanism; RobustRAG (Keyword) [61], which is designed for adversarial robustness; AstuteRAG [58], a framework designed to handle knowledge conflicts; and InstructRAG (with in-context learning) [60], which instructs LLMs to denoise contexts via rationales.
@@ 5.2 Evaluation Results for MIS
High Benign Performance.
The “benign” colum in Table 1 presents the results on benign data. Our MIS method consistently achieves high performance across all datasets and models. Compared specifically to RobustRAG (Keyword), which is also designed for robustness, MIS demonstrates significantly better benign performance across the board. Notably, on the long-form Biography generation task (Bio), MIS achieves high scores (e.g., 73 with Llama3.2-3B), markedly better than RobustRAG (Keyword) (56 with Llama3.2-3B). This highlights MIS’s ability to maintain utility for complex generation tasks, addressing a key limitation of previous robustness-focused methods. Our method also remarkably achieves comparable or sometimes superior accuracy compared to AstuteRAG and InstructRAG, even though these methods are designed for benign performance.
Robustness Against Adversarial Attacks.
Table 1 also shows the performance under prompt injection attack targeting either the highest-ranked (Position 1) or the lowest-ranked (Position 10) document. Our MIS method demonstrates substantial robustness. It significantly outperforms methods not explicitly designed for adversarial robustness such as Vanilla RAG, InstructRAG, and AstuteRAG. Compared to RobustRAG (Keyword), the other robustness-focused baseline, MIS also achieves better performance in general. Crucially, MIS retains its strong performance on the Bio long-form generation task even under attack (e.g. 71 @Pos 1 and 72.1 @ Pos 10 with Llama3.2-3B), whereas RobustRAG (Keyword) still struggles significantly on this task (53 @Pos 1 and 51.9 @Pos 10 with Llama3.2-3B). This underscores the advantage of MIS for robust long-form generation.
@@ 5.3 Evaluation Results for Sampling + MIS
In terms of benign performance, Sampling + MIS achieves strong utility across different models and datasets, competitive with or exceeding baselines like Vanilla RAG and InstructRAG. Notably, using GPT-4o-mini, Sampling + MIS consistently delivered the highest benign accuracy or LLM-Judge score across all four datasets compared to all baselines. Regarding robustness under prompt injection attack, our method shows significant resilience, particularly demonstrating the value of reliability awareness. Across models and datasets, Sampling + MIS almost always achieves the highest robust accuracy when the attack targets middle-ranked (Position 25) or low-ranked (Position 50) documents, scenarios where adversarial document corruption might be more feasible. When the attack targets the highest-ranked document (Position 1), the performance of Sampling + MIS is sometimes slightly lower than the best baseline in certain settings but remains competitive overall. These results indicate that the Sampling + MIS framework effectively scales the benefits of reliability-aware robustness to larger document sets, maintaining high utility while offering strong protection, especially against attacks targeting less reliable, lower-ranked information. In Appendix 5.5, we further demonstrate the reliability-awareness of our approach by plotting performance against attack positions.
5.4 Evaluation Results for Multi-Position Attacks

### v2

@@ 3.2 Robustness Analysis
Performance of Algorithm 1 depends on both NLI and MIS, and thus, of course, the robustness of our proposed system will depend on how NLI treats malicious documents. We assume that the NLI model has an error probability of at most $\epsilon_{1}$ when comparing two benign answers (i.e., it incorrectly labels them as contradictory), and an error probability of at most $\epsilon_{2}$ when comparing a benign answer and a malicious answer (i.e., it incorrectly fails to detect contradiction). Formally, for each pair of answers $(y_{i},y_{j})$ produced from documents $(x_{i},x_{j})$: if both $x_{i}$ and $x_{j}$ are benign, then $\text{NLI}(y_{i},y_{j})$ outputs “non-contradictory” with probability at least $1-\epsilon_{1}$; if exactly one of $x_{i},x_{j}$ is malicious, then $\text{NLI}(y_{i},y_{j})$ outputs “contradictory” with probability at least $1-\epsilon_{2}$. We place larger tolerance on $\epsilon_{2}$ because adversaries may craft malicious documents in such a way that induces larger NLI error rates. We make no assumption on NLI output when both are malicious.
Theorem 1.
@@
The assumption that NLI is able to find contradictions over LLM’s isolated answers is justified by the targeted attack we consider in our threat model, where an adversary aims to manipulate the model’s output to induce a specific outcome, such as inclusion of their own product in Google AI Overview. Thus, for the attack to be meaningful in this setting, the malicious document should diverge from the information in benign documents, promoting an alternative product that would not otherwise appear in the output. Therefore, even in the case of prompt injection, the injected content is crafted to induce such a targeted outcome, which ensures that the resulting answers will diverge from benign ones and can be detected as contradictions by NLI. The assumption of bounded corruption stems from the practical difficulty of manipulating many top-ranked documents in RAG-based search systems that rely on the strength of modern information retrieval systems against SEO Joachims et al. (2005); Ankalkoti (2017); Shahzad et al. (2020).
For the NLI model, we use DeBERTa-v3-large-mnli-fever-anli-ling-wanli Laurer et al. (2024) checkpoint, which achieves state-of-the-art 91.2% / 90.8% accuracy on the MNLI-m/mm test splits (benign) and 70.2% on the ANLI test set (adversarial). These numbers indicate both high everyday reliability and strong robustness to deliberately hard contradictions.
Theorem 1 (proved in Appendix B.1) provides theoretical validation that ReliabilityRAG via MIS is provably robust since the probability of selecting a malicious document vanishes as the number of retrieved documents $k$ grows large, provided the NLI error rates $\epsilon_{1},\epsilon_{2}$ and the number of malicious documents $k^{\prime}$ satisfy the specified conditions. To better understand the practical robustness implications in small-$k$ regimes, we present empirical results demonstrating practical robustness in such regimes in Appendix B.1.1. In addition, in Appendix B.2, we show that when $\epsilon_{1}=\epsilon_{2}=0$ (perfect NLI), the MIS is exactly the set of benign documents when $k^{\prime}<\frac{k}{2}$, i.e., we guarantee perfect robustness and utility.
@@ 4 Cardinal-Reliability Setting: Weighted Sample and Aggregate Framework
In addition, thus far, we have primarily focused on the ordinal-reliability setting. However, many retrieval systems provide, in addition to a rank, a non-negative reliability score $r(x_{i})$ for each retrieved document $x_{i}$ Elastic (2025). We can naturally convert these scores into normalized weights by setting $w(x_{i})=\frac{r(x_{i})}{\sum_{j=1}^{k}r(x_{j})}$, so a larger weight signifies greater trustworthiness. Cardinal reliability therefore makes the distribution of trust explicit: In the search scenario, depending on the query, one could encounter various weight distributions. For example, for queries with many reliable sources, such as encyclopedia-type queries, the weights among the top $k$ retrieved documents may be rather uniform; on the other hand, for niche queries, the highest ranked documents may carry high weights with a sharp drop-off for the lower ranked ones.
@@ 4.1 Weighted Sample and Aggregate Framework
Instantiation of $\mathcal{A}$ with MIS-based Document Selection. We can instantiate the aggregator $\mathcal{A}$ with MIS-based document selection (Stage 2 and 3 of Algorithm 1) by sending $a_{t}$’s as the isolated answers and contexts $S_{t}$’s as the documents. The ranking over the contexts in the instantiation is defined based on the rankings of the documents inside them. Since each context $\mathcal{S}_{t}$ is a tuple of documents, we can rank them lexicographically based on the ranks of the documents they contain. [^4] Again, this is one of many possible design choices. With this instantiation, we are able to control the running time of MIS via the choice of the number of sampling rounds $T$, as the MIS is now computed on a graph with $T$ vertices instead of $k$.
@@ 5.1 Experimental Setup
Datasets. We evaluate on three open-domain QA datasets: RealtimeQA (RQA) Kasai et al. (2023), Natural-Questions (NQ) Lee et al. (2019), TriviaQA (TQA) Joshi et al. (2017), and a long-form Biography generation dataset (Bio) Lebret et al. (2016). These datasets have been widely used to study the accuracy and robustness of RAG systems Xiang et al. (2024a); Wei et al. (2025); Wang et al. (2024a), making them well-suited benchmarks for evaluating our method. The detailed setup for datasets is presented in Appendix C.1. Due to limited space, we present results for TQA in Appendix C.2, which largely mirror the results for RQA and NQ.
LLMs and RAG Settings. We run experiments using three LLMs as the generators in our RAG pipelines: Mistral-7B-Instruct-v0.2 Jiang et al. (2023), Llama3.2-3B-Instruct Meta (2024), and GPT-4o-mini OpenAI (2024b). We set temperature to 0 for all experiments. When testing MIS, we use the top $k=10$ passages. For Sampling + MIS, we use the top $k=50$ documents, since one major motivation for the weighted sample and aggregate framework is scalability. We set context size $m=2$ and number of sampling rounds $T=20$. For the weights, we use the exponentially decaying weights and set $w(x_{i})\propto\gamma^{i-1}$, where $\gamma=0.9$. We present detailed ablation studies and discussions on the choice of parameters and their impact in Appendix C.6.
Evaluation Scenarios. We assess our algorithm’s performance and robustness under two distinct scenarios: a benign setting where no adversarial attack is performed and an adversarial attack setting. We consider the benign setting as defenses risk filtering out helpful content along with harmful attacks, reducing accuracy even when there is no attack Xiang et al. (2024a). For the adversarial setting, we simulate targeted corruption of one document at specific ranks: positions 1 (highest) and 10 (lowest) for $k=10$, and positions 1 (highest), 25 (middle), and 50 (lowest) for $k=50$. We also evaluate on multi-position attacks. Due to limited space, we present partial evaluation results for multi-position attacks in Section 5.4 and full experimental settings and results in Appendix C.5.
@@
Evaluation Metrics. We evaluate all methods based on their ability to generate accurate responses, both in benign conditions and under attack. We report GPT-4o-judged answer-accuracy for QA datasets and a GPT-4o judge score (0–100) for Bio. Due to limited space, we evaluate Attack Success Rate (ASR) in Appendix C.4.2.
Baselines. We compare our proposed reliability-aware methods against several baselines: Vanilla RAG, which concatenates all retrieved passages without any defense mechanism; RobustRAG (Keyword) Xiang et al. (2024a), which is designed for adversarial robustness; AstuteRAG Wang et al. (2024a), a framework designed to handle knowledge conflicts; and InstructRAG (with in-context learning) Wei et al. (2025), which instructs LLMs to denoise contexts via rationales.
@@ 5.2 Evaluation Results for MIS
High Benign Performance. The “benign” colum in Table 1 presents the results on benign data. Our MIS method consistently achieves high performance across all datasets and models. Compared specifically to RobustRAG (Keyword), which is also designed for robustness, MIS demonstrates significantly better benign performance across the board. Notably, on the long-form Biography generation task (Bio), MIS achieves high scores (e.g., 73 with Llama3.2-3B), markedly better than RobustRAG (Keyword) (56 with Llama3.2-3B). This highlights MIS’s ability to maintain utility for complex generation tasks, addressing a key limitation of previous robustness-focused methods. Our method also remarkably achieves comparable or sometimes superior accuracy compared to AstuteRAG and InstructRAG, even though these methods are designed for benign performance.
Robustness Against Adversarial Attacks. Table 1 also shows the performance under prompt injection attack targeting either the highest-ranked (Position 1) or the lowest-ranked (Position 10) document. Our MIS method demonstrates substantial robustness. It significantly outperforms methods not explicitly designed for adversarial robustness such as Vanilla RAG, InstructRAG, and AstuteRAG. Compared to RobustRAG (Keyword), the other robustness-focused baseline, MIS also achieves better performance in general. Crucially, MIS retains its strong performance on the Bio long-form generation task even under attack (e.g. 71 @Pos 1 and 72.1 @ Pos 10 with Llama3.2-3B), whereas RobustRAG (Keyword) still struggles significantly on this task (53 @Pos 1 and 51.9 @Pos 10 with Llama3.2-3B). This underscores the advantage of MIS for robust long-form generation.
@@ 5.3 Evaluation Results for Sampling + MIS
In terms of benign performance, Sampling + MIS achieves strong utility across different models and datasets, competitive with or exceeding baselines like Vanilla RAG and InstructRAG. Notably, using GPT-4o-mini, Sampling + MIS consistently delivered the highest benign accuracy or LLM-Judge score across all four datasets compared to all baselines. Regarding robustness under prompt injection attack, our method shows significant resilience, particularly demonstrating the value of reliability awareness. Across models and datasets, Sampling + MIS almost always achieves the highest robust accuracy when the attack targets middle-ranked (Position 25) or low-ranked (Position 50) documents, scenarios where adversarial document corruption might be more feasible. When the attack targets the highest-ranked document (Position 1), the performance of Sampling + MIS is sometimes slightly lower than the best baseline in certain settings but remains competitive overall. These results indicate that the Sampling + MIS framework effectively scales the benefits of reliability-aware robustness to larger document sets, maintaining high utility while offering strong protection, especially against attacks targeting less reliable, lower-ranked information. In Appendix C.4, we further demonstrate the reliability-awareness of our approach by plotting performance against attack positions.
5.4 Partial Evaluation Results for Multi-Position Attacks
Figure 2: Accuracy versus number of attacked documents on NQ
Due to limited space, we only present evaluation results for Sampling + MIS on NQ with $k=50$ retrieved documents here, and leave detailed experimental settings and results on MIS and other datasets for Appendix C.5. We compare with RobustRAG (Keyword) as a baseline. We craft a cleaned version of the dataset where all documents are relevant to allow for a clearer understanding of the impact of number of attacked documents, and attack a suffix of documents (e.g. documents at positions 46 - 50), echoing our reliability-aware setting. In Figure 2, we plot accuracy versus number of attacked documents. We see that, even as the attacker corrupts up to 40% of the passages, Sampling + MIS shows decent performance, whereas RobustRAG (Keyword) collapses much faster. Results on other datasets and for MIS with $k=10$ retrieved documents follow the same graceful-degradation pattern.
6 Conclusion
Acknowledgments and Disclosure of Funding
Appendix A Related Works
Adversarial Attacks Against RAG. The standard RAG pipeline involves retrieving relevant documents, optionally re-ranking them, and feeding them to the LLM for generation. However, this reliance on external data creates vulnerabilities. Early works studied misinformation attacks against QA models. Recent attacks specifically target LLM-powered RAG, evolving rapidly. Corpus poisoning or retrieval corruption involves injecting malicious content into the knowledge base; examples include PoisonedRAG Zou et al. (2025), Topic-FlipRAG Gong et al. (2025), MM-PoisonRAG Ha et al. (2025). Prompt injection, embedding malicious instructions in retrieved data, remains a top threat. Other methods include low-level perturbations such as typos (e.g., GARAG Cho et al. (2024)) and increasingly, data extraction attacks aiming to steal information from the RAG database using optimization techniques (MARAGE Hu et al. (2025)), backdoors implanted during fine-tuning, or automated agent-based methods (RAG-Thief Jiang et al. (2024)). Some recent works demonstrate the manipulability of LLM preferences concerning products, a significant issue as LLMs are increasingly used for product recommendations. For instance, Kumar and Lakkaraju (2024) show that inserting a “strategic text sequence” into product metadata can improve its ranking in LLM outputs, while Nestaas et al. (2024) found that adversarial content on product webpages can alter its own or competitors’ rankings in LLM recommendations.
Robust RAG Frameworks. Several frameworks aim to improve RAG robustness, addressing resilience against noise or defense against adversarial attacks:
  • RobustRAG: Xiang et al. (2024a) Employs an “isolate-then-aggregate” strategy (using keyword and decoding aggregation) for provable robustness against retrieval corruption. However, their methods have limited scalability to long-form generation. The keyword-based voting, for instance, struggles when answers are lengthy or complex, because it generates the ultimate answer based only on a few keywords and necessarily suffers from significant information loss.
  • InstructRAG: Wei et al. (2025) Teaches LLMs to denoise context via self-synthesized rationales, improving robustness to noisy retrievals without extra supervision. However, as shown in our evaluations and Zhou et al. (2025), it is not robust against simple adversarial attacks such as prompt injection.
  • AstuteRAG: Wang et al. (2024a) Addresses imperfect retrieval and internal/external knowledge conflicts by eliciting internal knowledge, consolidating sources, and selecting the most reliable answer. It doesn’t explicitly use initial retrieval rank and has been primarily evaluated on short-form QA. Similar to InstructRAG, as shown in our evaluations and Zhou et al. (2025), it is not robust against simple adversarial attacks such as prompt injection.
  • TrustRAG: Zhou et al. (2025) Defends against corpus poisoning using $k$-means clustering to filter suspicious documents and LLM self-assessment to resolve conflicts. However, it makes the unrealistic assumption that malicious documents form a separate cluster in the embedding space. This assumption is particularly challenging given that precisely controlling such representations is difficult, and adversaries would naturally strive to craft malicious content to be semantically similar to benign documents to evade detection Zou et al. (2025). In contrast, our assumption centers on the established capabilities of modern NLI models to discern contradictions, even if imperfectly and in adversarial settings, which is a more tenable premise in an adversarial context where attackers prioritize stealth over creating easily detectable patterns.
Document Reweighting, Selection and Filtering. A complementary line of work seeks to filter or reweigh retrieved passages before any generation occurs, ensuring that the context presented to the language model is already relevant, self-consistent, and trustworthy.
  • Self-RAG: Asai et al. (2024) let the LLM emit “reflection” tokens that mark useless documents and trigger additional retrieval. However, it ignores source-level reliability and offers no provable guarantees.
  • Chain-of-Note: Yu et al. (2024) has the LLM sequentially read each retrieved document and write a brief “note” assessing its relevance before attempting an answer. If a document is deemed unhelpful in these notes, it can be effectively filtered out and not used in the final answer reasoning.
  • CRAG: Yan et al. (2024) introduced a retrieval evaluator that scans the retrieved set and predicts whether the question is answerable with the given documents. If deemed unanswerable, the system can abstain or retrieve from a broader source, rather than force a guess.
  • CrAM: Deng et al. (2025) scores each passage with an external credibility estimator and down-weights low-credibility tokens in the LLM’s attention.
  • RA-RAG: Hwang et al. (2025) Explicitly models source reliability using iterative offline estimation and weighted majority voting aggregation. However, it focuses primarily on reliability estimation instead of aggregation. It has also not been shown to be robust against adversarial attacks.
Sampling-Based Methods for Robustness. Sampling strategies have long been used to bolster model robustness under adversarial or noisy data conditions. A classic example is Random Sample Consensus (RANSAC) Bathula (2023), which fits models on randomly sampled data subsets to ignore outliers and find a consensus solution. Modern defenses introduce stochasticity to blunt adversarial attacks. For example, Xie et al. (2018) proposes adding random transformations at inference time, which demonstrate effective mitigation of adversarial image perturbations without requiring specialized training.
Recent works have also leveraged sampling for provable robustness. Cohen et al. (2019) introduces randomized smoothing, which converts any classifier into a certifiably robust one by adding Gaussian noise to inputs and predicting via majority vote. Ben-Eliezer and Yogev (2020) show that a standard reservoir sampling can be made robust to an adversarial input stream by increasing the sample size. Their analysis illustrates how strategic resampling can ensure a representative sample despite an attacker’s attempts to corrupt the data.
Reliability Signals in Ranking and Recommendation Systems. Reliability signals are also useful in protecting ranking and recommendation systems against spam and manipulation. Gyongyi et al. (2004) propagates trust from a small seed of human-vetted pages through the web graph, sharply demoting sites that lie far from trusted regions and reducing the impact of link-spam on search results. Massa and Avesani (2007) shows that embedding explicit user-trust scores into collaborative filtering not only boosts accuracy but also curbs profile-injection attacks, demonstrating the defensive value of trust-weighted aggregation.
Appendix B Theoretical Analysis and Proofs
B.1 Analysis of MIS Robustness with Imperfect NLI
In this section, we present the proof of Theorem 1. Note that for the proof, we assume that the errors for each edge occur independently.
Theorem (Theorem 1 restated).
Proof of Theorem 1.
Thus, by union bound, the probability that a malicious document ends up in the maximum independent set is at most $\Pr[\textsc{Bad}_{1}]+\Pr[\textsc{Bad}_{2}]\leq e^{-O(k)}$, so the probability that the maximum independent set does not contain any malicious document is at least $1-e^{-O(k)}$, finishing the proof. ∎
B.1.1 Simulations on Small $k$
(a) $k=10$, $\epsilon_{1}=0.05$, $\epsilon_{2}=0.2$
(b) $k=10$, $\epsilon_{1}=0.05$, $\epsilon_{2}=0.4$
(c) $k=20$, $\epsilon_{1}=0.05$, $\epsilon_{2}=0.2$
(d) $k=20$, $\epsilon_{1}=0.05$, $\epsilon_{2}=0.4$
B.2 Analysis of MIS Robustness with Perfect NLI
Theorem 2.
Proof.
  1. 1.
  2. 2.
Thus, the set of benign documents $B$ is the maximum independent set. ∎
B.3 Robustness Guarantee for Weighted Sampling
Each of the $T$ rounds of sampling represents an independent Bernoulli trial with success probability $p_{\mathrm{clean}}$, where success means drawing a clean context. Let $C$ be the total number of clean contexts generated across the $T$ rounds. $C$ follows a binomial distribution: $C\sim\operatorname{Binomial}(T,p_{\mathrm{clean}})$. In Theorem 3, we present the robustness guarantee of Algorithm 2.
Theorem 3.
$T\geq\frac{1}{2(p_{\mathrm{clean}}-(1-\alpha))^{2}}\log\frac{1}{\delta},$
Proof of Theorem 3.
By Hoeffding’s inequality, we have
Thus, the probability of success is
With $T\geq\frac{1}{2(p_{\mathrm{clean}}-(1-\alpha))^{2}}\log\frac{1}{\delta}$, we have $\Pr[C\geq T(1-\alpha)]\geq 1-\delta$, so Algorithm 2 instantiated with $\mathcal{A}$ is $(1-\delta)\lambda$-robust. ∎
Concrete Instantiation. Take $\alpha=\frac{1}{2}$, $m=2$ and $\eta=0.1$. Then $p_{\mathrm{clean}}=(1-0.1)^{2}\approx 0.81$. Since $p_{\mathrm{clean}}>1/2$, the condition for the Hoeffding bound is met. With $T=20$ we have a failure probability bound of: $\delta=\exp\left(-2\times 20\times(0.81-0.5)^{2}\right)\approx 0.0214.$ Thus, with these parameters, the algorithm returns a robust answer with probability at least 10.0214=0.97861-0.0214=0.9786, or $97.86\%$.
Appendix C Additional Experimental Setup and Evaluation Results
C.1 Detailed Experimental Setup
Datasets. We evaluate our methods on both short-answer open-domain question answering (QA) and long-form text generation tasks. For QA, we use RealtimeQA (RQA) Kasai et al. (2023), Natural Questions (NQ) Lee et al. (2019), and TriviaQA (TQA) Joshi et al. (2017). For long-form generation, we utilize the Biography generation dataset (Bio) Lebret et al. (2016). We use 100 queries from RQA dataset, randomly draw 500 queries from each of NQ dataset and TQA dataset, and 50 queries from Bio dataset. For RQA, we use the 100 queries provided by Xiang et al. (2024a). This differs from the 500 queries sampled for NQ and TQA because the RobustRAG work, which serves as our primary baseline for comparison, only included this specific set of 100 RQA queries in their Git Repo Xiang et al. (2024b). This circumvents issues arising from RQA’s real-time nature, as the dataset has not been actively updated recently (latest public data points appear to be from 2023), making it problematic to use current search results for its potentially outdated questions. For each query, we retrieve relevant passages using Google Search. Since we initially retrieve only the search result snippets displayed on the first page, crucial information is often truncated (indicated by "…"). Consequently, the initial ranking provided by the search engine may not accurately reflect the relevance or quality of the snippet content. To address this, we re-rank the retrieved passages for each query using the mxbai-rerank-large-v2 model Lee et al. (2025). This re-ranking step is a common practice in modern RAG pipelines to enhance context quality and can be performed efficiently.
Evaluation Metrics. The detailed evaluation metrics we use are:
  • For QA tasks (RQA, NQ, TQA): We assess correctness by comparing the generated answer $r$ against the gold answer $g$. GPT-4o serves as an LLM judge to classify the answer as correct or incorrect based on Zheng et al. (2023). The reported metric is Accuracy %, representing the percentage of correctly answered queries.
  • For the long-form Bio generation task: We evaluate the quality of the generated biography $r$ following a multi-aspect LLM-as-a-judge rubric similar to Liu et al. (2023). First, a reference (gold) response $g$ is generated by prompting GPT-4o with the full Wikipedia document of the target person. Subsequently, GPT-4o serves as an LLM judge to compare $r$ against $g$, providing individual scores from 0 to 10 for three distinct criteria: (i) factual accuracy, (ii) relevance and recall, and (iii) coherence and structure. The exact prompt template for grading is presented in Appendix E.2. For each query, these three scores are averaged and scaled to 100. The reported metric, the LLM-Judge Score, is the average of these final per-query scores across all queries evaluated in the dataset.
C.2 Evaluation Results for TQA under Prompt Injection Attack
C.3 Evaluation Results for Corpus Poisoning Attack
MIS ($k=10$). The MIS method demonstrates strong robustness against poisoning attacks, generally outperforming all other baselines. It mostly maintains better performance, especially on the long-form Bio task, compared to RobustRAG (Keyword). Rank-awareness is evident, with MIS typically performing better when the attack targets Position 10 versus Position 1.
Sampling + MIS ($k=50$). Similarly, Sampling + MIS shows good resilience. It achieves high empirical accuracy under attacks, particularly when attacks target mid- or lower-ranked documents (positions 25 and 50), reinforcing the effectiveness of reliability-aware sampling framework against poisoning in larger document sets.
C.4 Reliability-Awareness of Our Methods
C.4.1 Accuracy Versus Attack Position
C.4.2 Attack Success Rate Versus Attack Position
C.5 Full Evaluation Results for Multi-Position Attacks